# QuMAB: Query-based Multi-annotator Behavior Pattern Learning

### v1

@@ Abstract.
Multi-annotator learning traditionally aggregates diverse annotations to approximate a single “ground truth”, treating disagreements as noise. However, this paradigm faces fundamental challenges: subjective tasks often lack absolute ground truth, and sparse annotation coverage makes aggregation statistically unreliable. We introduce a paradigm shift from sample-wise aggregation to annotator-wise behavior modeling. By treating annotator disagreements as valuable information rather than noise, modeling annotator-specific behavior patterns can reconstruct unlabeled data to reduce annotation cost, enhance aggregation reliability, and explain annotator decision behavior. To this end, we propose QuMATL (Query-based Multi-Annotator Behavior Pattern Learning), which uses lightweight queries to model individual annotators while capturing inter-annotator correlations as implicit regularization, preventing overfitting to sparse individual data while maintaining individualization and improving generalization, with a visualization of annotator focus regions offering an explainable analysis of behavior understanding. We contribute two large-scale datasets with dense per-annotator labels: STREET (4,300 labels/annotator) and AMER (average 3,118 labels/annotator), the first multimodal multi-annotator dataset.
@@
ccs: Computing methodologies Multi-task learning
@@ 1. Introduction
Our key insight is that annotators, despite their individual differences, often share behavioral structures. By capturing inter-annotator correlations, we can leverage the collective patterns as implicit regularization, constraining individual models from overfitting while preserving unique characteristics. To this end, we propose a novel query-based architecture QuMATL (Query-based Multi-Annotator Behavior Pattern Learning), hypothesize that annotator judgment differences arise from their varying degrees of focus on different regions of the input content (e.g., focusing on different image patches). Each annotator is represented by a learnable query that interacts with input features via cross-attention to effectively model individual behavior patterns. Lightweight query significantly reduces computational cost compared to separate conventional models. Crucially, all annotator queries also interact through shared self-attention to capture inter-annotator correlations as a form of implicit structural regularization. This constrains inter-annotator representations to follow similarity patterns derived from annotations, preventing individual representations from drifting too far from the group and promoting mutual enhancement. This mechanism prevents overfitting to sparse individual data while maintaining individualization, improving generalization and robustness in individual annotator modeling, particularly under sparse annotations. Additionally, the cross-attention weights provide a visualization of annotator focus regions, offering an explainable analysis of behavior understanding.
@@
  • A novel query-based architecture: We propose QuMATL, which uses lightweight queries to model individual annotators while capturing inter-annotator correlations as implicit regularization, preventing overfitting to sparse individual data while maintaining individualization and improving generalization, with a visualization of annotator focus regions offering an explainable analysis of behavior understanding.
@@
MUSIC (Rodrigues et al., 2013a) music genre classification audio 2$\sim$368
@@
AMER (Ours) video emotion recognition audio, video, text 970$\sim$5,202
@@ 2.1. Multi-annotator Behavior Modeling Paradigm
To the best of our knowledge, the multi-annotator behavior modeling paradigm problem has not yet been investigated. Traditional multi-annotator learning focuses on estimating consensus or ground-truth labels from multiple noisy annotations. These include early probabilistic models (Dawid and Skene, 1979), EM algorithms (Whitehill et al., 2009), Gaussian models (Rodrigues et al., 2014), and biased estimation (Welinder et al., 2010). Tanno et al. (Tanno et al., 2019b) proposed modeling annotator confusion matrices as learnable parameters in neural networks. Cao et al. (Cao et al., 2019) introduced max-MIG to learn from multiple annotators. NEAL (Chen et al., 2023) employs neural expectation-maximization to jointly learn annotator expertise and true labels. Later methods used probabilistic frameworks to aggregate multiple annotations into a consensus or ground-truth label by confusion matrix (Tanno et al., 2019a), agreement distribution (Wang et al., 2023), and Gaussian distributions (Liao et al., 2024). This sample-wise aggregation paradigm often treats annotator disagreements as noise to be averaged away rather than valuable information (Yan et al., 2014; Jinadu et al., 2023). In contrast, our introduced annotator-wise modeling paradigm treats annotator disagreements as valuable information for modeling annotator-specific behavior patterns, enhancing aggregation reliability, and explaining annotator behavior.
@@ 2.3. Multi-annotator Datasets
Figure 2. QuMATL for image. A frozen pre-trained image encoder extracts features, which interact with annotator-specific learnable queries in the Image Q-Former through cross-attention, producing annotator-specific features for classification. Different annotators’ cross-attention weights reflect the differences in the image patches they focus on and support visual interpretation.
@@ 3. Dataset Construction
(1) STREET is an urban perception dataset with multi-annotators, which contains $4,300$ high-resolution images covering various urban elements, such as streets, public spaces, and infrastructure. The images were captured during a series of city strolling surveys, which aim to analyze emotions in relation to various factors associated with the city. The surveys were conducted by an organization to which one of our co-authors belongs. Voluntary participants walked around their own familiar city, took photos of various factors that may affect their subjective feelings (i.e., happiness/health), and assigned labels related to these feelings to each image (though we do not use these labels but ones assigned by crowd workers in our experiments). Thirteen survey sessions were conducted in five different cities (three urban areas and two suburban areas). A total of 327 participants, ranging in age from their 10s to 60s, took part in the survey. Each session lasted about one hour, with each participant taking an average of 12.6 photos. We outsourced the annotation process to a company, which selected $10$ annotators with balanced age, gender, and location diversity on a platform similar to Amazon Mechanical Turk, assessing five perception dimensions: happiness, healthiness, safety, liveliness, and orderliness, using a $6$-point scale ($-3$ to $+3$). Each annotator spent approximately three weeks on their annotations. This multi-annotator dataset provides comprehensive human perception data for urban environments, enabling the quantitative analysis of environmental features and their emotional impact.
@@ 4. Methodology
We propose QuMATL, a query-based architecture designed to model the behavior patterns of individual annotators. We take the image classification task to illustrate our image-specific architecture for image inputs from the STREET dataset, which consists of a frozen image encoder, annotator-specific learnable queries, an image Q-Former, and annotator-wise classifiers, as shown in Figure 2. For video inputs from the AMER dataset, we describe a video-specific architecture in the supplementary material.
@@ 4.1. Loss Function
Finally, as shown in Figure 2, the total training loss $\mathcal{L}_{\text{total}}$ for the proposed multi-annotator classification model, QuMATL, is defined as the sum of individual cross-entropy losses for each annotator:
@@ 5. Experiment
We conduct extensive experiments to evaluate our QuMATL, including modeling individual annotators’ behavior patterns, assessing their utility for consensus prediction, testing applicability under sparse annotations, and complementing with qualitative visualization analysis. We compare against three representative baselines: D-LEMA (Mirikharaji et al., 2021), an ensemble-based multi-annotator learner; PADL (Liao et al., 2024), which fits Gaussian distributions for each annotator; and MaDL (Herde et al., 2023), which models annotator-specific confusion matrices. To ensure our model captures annotator-specific patterns rather than shared encoder features, we also include a Base variant with only the encoder and classifiers. Results are reported on two novel datasets. Accuracy and $F_{1}$ score (Zhan et al., 2019) are used as evaluation metrics. Note that additional experiments, including model efficiency, a faithfulness-oriented interpretability analysis, extended results, and further discussion, are provided in the supplementary material.
@@ 5.1. Implementation Details
Our image-specific model pipeline (Figure 2) uses ViT-G/14 from EVA-CLIP (Sun et al., 2023) as the encoder, with Image Q-Former initialized from InstructBLIP (Dai et al., 2023) (Frame Q-Former is the same in a video-specific pipeline from supplementary material, where Video Q-Former is initialized from Video-LLaMA (Zhang et al., 2023b)). Input images and video frames are resized to 224$\times$224 and normalized. The number of query tokens is set equal to the number of annotators, and each annotator’s classifier model uses an MLP. We train the model using the AdamW optimizer with an initial learning rate of 1e-4, a weight decay of 0.01, and gradient clipping with a maximum norm of 1.0. A linear warmup strategy is applied for the first 20% steps followed by cosine learning rate decay. The model is trained for up to 200 epochs with early stopping (patience = 25) to avoid overfitting. Training is conducted using distributed data parallelism (DDP) on four NVIDIA V100 GPUs.
@@ 6. Conclusion
This paper introduced a paradigm shift in multi-annotator learning from sample-wise aggregation to annotator-wise behavior modeling, and proposed a lightweight query-based architecture (QuMATL) to effectively model individual annotator behavior patterns. We contributed the STREET and AMER datasets with dense per-annotator labels for longitudinal annotator behavior understanding. Experiments demonstrate the superiority of our method in modeling individual annotators’ behavior patterns, their utility for consensus prediction, and applicability under sparse annotations. This work presents a novel view on recognizing and understanding the challenges in the multi-annotator learning field.
Figure 5. QuMATL architecture for video-specific pipeline. A frozen pre-trained encoder extracts frame features, which are compressed through the Frame Q-Former. With added frame position embedding, these interact with $1,\dots,n$ annotator-specific learnable queries in the Video Q-Former through cross-attention, producing annotator-specific features for classification. Different annotators’ cross-attention weights represent different annotator tendencies and provide visualization for analysis.
@@ 7.1. Overview
In this supplementary, we first illustrate the detailed video-specific architecture in Section 7.2. We provide the analysis of model efficiency in Section 7.3. Also, the extended discussion is provided in Section 7.4. Finally, we show the additional experimental results in Section 7.5.
@@ 7.2. Video-specific Architecture
Different from the image input pipeline in the main paper, Figure 5 presents the video input pipeline specifically designed for the AMER, etc. video datasets. Specifically, given a video input $V\in\mathbb{R}^{T\times H\times W\times 3}$, a frozen pre-trained image encoder (Sun et al., 2023) first extracts frame features, which are then further compressed by the image Q-Former using a specific number of compression queries, typically 32, to alleviate subsequent computational costs (Li et al., 2023). The frame position embedding is then added and input to the video Q-Former. Subsequently, we define the unique tendency for annotator $A_{k}$ ($k=1,\dots,n$) through each learnable query in a Q-Former, then all annotator-specific queries in cross-attention simultaneously interact with input features to capture their diverse tendencies and obtain annotator-specific feature. Finally, the specific annotator features are mapped through a fully connected layer to an appropriate feature dimension and then connected to each annotator’s corresponding classifier to output classification results.
For visualization, the focus points to video input are based on all frames of the video. Different annotators’ varying levels of focus on different frames indicate their tendency differences: annotator 1 shows more focus on the final frames, while annotator n focuses more on the middle frames. This dataset is video emotion classification. Analyzing the annotation and visualization results, we see that annotators labeled emotions “anger” and “worried”. The differences in the frame segments they focused on might partially contribute to the variations in their judgments.

### v2

# QuMAB: Query-based Multi-Annotator Behavior Modeling with Reliability under Sparse Labels

@@ Abstract.
Multi-annotator learning traditionally aggregates diverse annotations to approximate a single “ground truth”, treating disagreements as noise. However, this paradigm faces fundamental challenges: subjective tasks often lack absolute ground truth, and sparse annotation coverage makes aggregation statistically unreliable. We introduce a paradigm shift from sample-wise aggregation to annotator-wise behavior modeling. By treating annotator disagreements as valuable information rather than noise, modeling annotator-specific behavior patterns can reconstruct unlabeled data to reduce annotation cost, enhance aggregation reliability, and explain annotator decision behavior. To this end, we propose QuMAB (Query-based Multi-Annotator Behavior Pattern Learning), which uses lightweight queries to model individual annotators while capturing inter-annotator correlations as implicit regularization, preventing overfitting to sparse individual data while maintaining individualization and improving generalization, with a visualization of annotator focus regions offering an explainable analysis of behavior understanding. We contribute two large-scale datasets with dense per-annotator labels: STREET (4,300 labels/annotator) and AMER (average 3,118 labels/annotator), the first multimodal multi-annotator dataset. Extensive experiments demonstrate the superiority of our QuMAB in modeling individual annotators’ behavior patterns, their utility for consensus prediction, and applicability under sparse annotations.
@@
$c$s: Computing methodologies Multi-task learning
@@ 1. Introduction
Our key insight is that annotators, despite their individual differences, often share behavioral structures. By capturing inter-annotator correlations, we can leverage the collective patterns as implicit regularization, constraining individual models from overfitting while preserving unique characteristics. To this end, we propose a novel query-based architecture QuMAB (Query-based Multi-Annotator Behavior Pattern Learning), hypothesize that annotator judgment differences arise from their varying degrees of focus on different regions of the input content (e.g., focusing on different image patches). Each annotator is represented by a learnable query that interacts with input features via cross-attention to effectively model individual behavior patterns. Lightweight query significantly reduces computational cost compared to separate conventional models. Crucially, all annotator queries also interact through shared self-attention to capture inter-annotator correlations as a form of implicit structural regularization. This constrains inter-annotator representations to follow similarity patterns derived from annotations, preventing individual representations from drifting too far from the group and promoting mutual enhancement. This mechanism prevents overfitting to sparse individual data while maintaining individualization, improving generalization and robustness in individual annotator modeling, particularly under sparse annotations. Additionally, the cross-attention weights provide a visualization of annotator focus regions, offering an explainable analysis of behavior understanding.
@@
  • A novel query-based architecture: We propose QuMAB, which uses lightweight queries to model individual annotators while capturing inter-annotator correlations as implicit regularization, preventing overfitting to sparse individual data while maintaining individualization and improving generalization, with a visualization of annotator focus regions offering an explainable analysis of behavior understanding.
@@
[…]
@@
[…]
@@ 2.1. Multi-annotator Behavior Modeling Paradigm
To the best of our knowledge, the multi-annotator behavior modeling paradigm problem has not yet been investigated. Traditional multi-annotator learning focuses on estimating consensus or ground-truth labels from multiple noisy annotations. These include early probabilistic models (Dawid and Skene, 1979), EM algorithms (Whitehill et al., 2009), Gaussian models (Rodrigues et al., 2014), and biased estimation (Welinder et al., 2010). Tanno et al. (Tanno et al., 2019b) proposed modeling annotator confusion matrices as learnable parameters in neural networks. Cao et al. (Cao et al., 2019) introduced max-MIG to learn from multiple annotators. SimLabel (Zhang et al., 2025) addresses the practical challenge of missing labels in multi-annotator scenarios. NEAL (Chen et al., 2023) employs neural expectation-maximization to jointly learn annotator expertise and true labels. Later methods used probabilistic frameworks to aggregate multiple annotations into a consensus or ground-truth label by confusion matrix (Tanno et al., 2019a), agreement distribution (Wang et al., 2023), and Gaussian distributions (Liao et al., 2024). This sample-wise aggregation paradigm often treats annotator disagreements as noise to be averaged away rather than valuable information (Yan et al., 2014; Jinadu et al., 2023). In contrast, our introduced annotator-wise modeling paradigm treats annotator disagreements as valuable information for modeling annotator-specific behavior patterns, enhancing aggregation reliability, and explaining annotator behavior.
@@ 2.3. Multi-annotator Datasets
Figure 2. QuMAB for image. A frozen pre-trained image encoder extracts features, which interact with annotator-specific learnable queries in the Image Q-Former through cross-attention, producing annotator-specific features for classification. Different annotators’ cross-attention weights reflect the differences in the image patches they focus on and support visual interpretation.
@@ 3. Dataset Construction
(1) STREET is an urban perception dataset with multi-annotators, which contains $4,300$ high-resolution images covering various urban elements, such as streets, public spaces, and infrastructure. The images were captured during a series of city strolling surveys, which aim to analyze emotions in relation to various factors associated with the city. The surveys were conducted by an organization to which one of our co-authors belongs. Voluntary participants walked around their own familiar city, took photos of various factors that may affect their subjective feelings (i.e., happiness/health), and assigned labels related to these feelings to each image (though we do not use these labels but ones assigned by crowd workers in our experiments). Thirteen survey sessions were conducted in five different cities (three urban areas and two suburban areas). A total of 327 participants, ranging in age from their 10s to 60s, took part in the survey. Each session lasted about one hour, with each participant taking an average of 12.6 photos. We outsourced the annotation process to a company, which selected $10$ annotators with balanced age, gender, and location diversity on a platform similar to Amazon Mechanical Turk, assessing five perception dimensions: happiness, healthiness, safety, liveliness, and orderliness, using a $6$-point scale (3-3 to +3+3). Each annotator spent approximately three weeks on their annotations. This multi-annotator dataset provides comprehensive human perception data for urban environments, enabling the quantitative analysis of environmental features and their emotional impact.
@@ 4. Methodology
We propose QuMAB, a query-based architecture designed to model the behavior patterns of individual annotators. We take the image classification task to illustrate our image-specific architecture for image inputs from the STREET dataset, which consists of a frozen image encoder, annotator-specific learnable queries, an image Q-Former, and annotator-wise classifiers, as shown in Figure 2. For video inputs from the AMER dataset, we describe a video-specific architecture in the supplementary material.
@@ 4.1. Loss Function
Finally, as shown in Figure 2, the total training loss $\mathcal{L}_{\text{total}}$ for the proposed multi-annotator classification model, QuMAB, is defined as the sum of individual cross-entropy losses for each annotator:
@@ 5. Experiment
We conduct extensive experiments to evaluate our QuMAB, including modeling individual annotators’ behavior patterns, assessing their utility for consensus prediction, testing applicability under sparse annotations, and complementing with qualitative visualization analysis. We compare against three representative baselines: D-LEMA (Mirikharaji et al., 2021), an ensemble-based multi-annotator learner; PADL (Liao et al., 2024), which fits Gaussian distributions for each annotator; and MaDL (Herde et al., 2023), which models annotator-specific confusion matrices. To ensure our model captures annotator-specific patterns rather than shared encoder features, we also include a Base variant with only the encoder and classifiers. We evaluate on two multi-annotator datasets: AMER (video-based emotion recognition with 13 annotators) and STREET (urban image impression assessment with 10 annotators across five dimensions: Happiness, Healthiness, Safety, Liveliness, and Orderliness). The AMER dataset’s complexity in capturing temporal emotion dynamics aligns with recent advances in time-sensitive emotion recognition (Zhang, 2024b; Zhang et al., 2024). These datasets provide dense, diverse annotations crucial for modeling annotator-specific behavior patterns. Accuracy and $F_{1}$ score (Zhan et al., 2019) are used as evaluation metrics. Note that additional experiments, including model efficiency, a faithfulness-oriented interpretability analysis, extended results, and further discussion, are provided in the supplementary material.
@@ 5.1. Implementation Details
Our image-specific model pipeline (Figure 2) uses ViT-G/14 from EVA-CLIP (Sun et al., 2023) as the encoder, with Image Q-Former initialized from InstructBLIP (Dai et al., 2023) (Frame Q-Former is the same in a video-specific pipeline from supplementary material, where Video Q-Former is initialized from Video-LLaMA (Zhang et al., 2023b)). Input images and video frames are resized to $\times224$ and normalized. The number of query tokens is set equal to the number of annotators, and each annotator’s classifier model uses an MLP. We train the model using the AdamW optimizer with an initial learning rate of 1e-4, a weight decay of 0.01, and gradient clipping with a maximum norm of 1.0. A linear warmup strategy is applied for the first 20% steps followed by cosine learning rate decay. The model is trained for up to 200 epochs with early stopping (patience = 25) to avoid overfitting. Training is conducted using distributed data parallelism (DDP) on four NVIDIA V100 GPUs.
@@ 6. Conclusion
This paper introduced a paradigm shift in multi-annotator learning from sample-wise aggregation to annotator-wise behavior modeling, and proposed a lightweight query-based architecture (QuMAB) to effectively model individual annotator behavior patterns. We contributed the STREET and AMER datasets with dense per-annotator labels for longitudinal annotator behavior understanding. Experiments demonstrate the superiority of our method in modeling individual annotators’ behavior patterns, their utility for consensus prediction, and applicability under sparse annotations. This work presents a novel view on recognizing and understanding the challenges in the multi-annotator learning field.
Figure 5. An illustration of paradigm shift from sample-wise aggregation to annotator-wise behavior modeling. (a) Existing methods overlook individual annotator information, potentially leading to suboptimal or biased consensus aggregation on real-world annotation matrices, where each annotator labels only a small subset of samples with disjoint coverage. (b) Our method captures each annotator’s longitudinal behavior patterns across their labeled samples, improving consensus aggregation reliability by leveraging reconstructed unlabeled data.
Figure 6. QuMAB architecture for video-specific pipeline. A frozen pre-trained encoder extracts frame features, which are compressed through the Frame Q-Former. With added frame position embedding, these interact with $1,\dots,n$ annotator-specific learnable queries in the Video Q-Former through cross-attention, producing annotator-specific features for classification. Different annotators’ cross-attention weights represent different annotator tendencies and provide visualization for analysis.
@@ 7.1. Overview
First, we provide an empirical analysis, elaborating on our first contribution in the paper, i.e., practical value of our paradigm shift from sample-wise aggregation to annotator-wise modeling, further clarifying our motivation and practical contribution. Subsequently, we illustrate the detailed video-specific architecture. Additionally, we present supplementary experimental analysis. The analysis of model efficiency and the discussion are given. Finally, the discussion of limitation and future work is provided.
Table 6. Aggregation accuracy under simulated missing labels in the test set. Modeling annotators before majority voting (S-Ours-MV) outperforms direct voting (S-MV) and the best-performing baseline (S-PADL-MV), validating the benefit of annotator-wise modeling for reliable aggregation.
Method S-Ha S-He S-Sa S-Li S-Or AMER
20% missing rate
S-MV 0.97 0.97 0.96 0.96 0.96 0.85
S-PADL-MV 0.97 0.98 0.97 0.96 0.97 0.86
S-Ours-MV 0.98 0.98 0.97 0.98 0.97 0.89
30% missing rate
S-MV 0.94 0.93 0.92 0.93 0.92 0.85
S-PADL-MV 0.95 0.95 0.95 0.94 0.94 0.87
S-Ours-MV 0.97 0.96 0.95 0.96 0.95 0.92
40% missing rate
S-MV 0.92 0.92 0.89 0.91 0.90 0.81
S-PADL-MV 0.93 0.93 0.91 0.92 0.91 0.83
S-Ours-MV 0.95 0.95 0.93 0.94 0.93 0.88
Table 7. Label statistics and missing rates of the AMER dataset with 13 annotators $A_{i},i\in\{1,\cdots,5\}$. For each annotator, the number of labeled samples, the corresponding missing rate (%), as well as the average data (Average), are provided.
Perspective $A_{1}$ $A_{2}$ $A_{3}$ $A_{4}$ $A_{5}$ $A_{6}$ $A_{7}$ $A_{8}$ $A_{9}$ $A_{10}$ $A_{11}$ $A_{12}$ $A_{13}$ Average
Sample size 1096 1031 1022 1036 1012 970 1064 1049 1060 1062 5187 5197 5202 1999.1
Missing rates (%) 79.0 80.2 80.4 80.1 80.6 81.4 79.6 79.9 79.6 79.6 0.4 0.2 0.1 69.6
7.2. Empirical Validation: Annotator Modeling Before Aggregation for Reliability
We present empirical validation and analysis to concretize our first contribution—the paradigm shift from sample-wise aggregation to annotator-wise behavior modeling—further clarifying our motivation and practical significance.
Importance of Annotator Modeling for Aggregation Reliability. As illustrated in Figure 5, traditional aggregation strategies operate along the sample axis without modeling longitudinal annotator behavior. In real-world sparse and non-overlapping scenarios, this leads to suboptimal or biased consensus aggregation: sample $S_{1}$ may be aggregated from annotators $A_{2},A_{3},A_{4}$, while sample $S_{2}$ from $A_{1},A_{3},A_{4}$—sometimes even entirely disjoint groups. Therefore, we hypothesize that modeling annotators individually allows for better reconstruction of their missing labels, thereby enabling more consistent and trustworthy aggregation across all samples (e.g., all $S_{1}$$S_{N}$ aggregated from $A_{1},A_{2},A_{3},A_{4}$) given well-trained annotator models.
Validation: Direct Aggregation vs. Aggregation after Annotator-wise Modeling. To validate this, as shown in Table 6, we simulate missingness on the test set by masking {20%, 30%, 40%} of annotations and compare: (1) S-MV: Direct sample-wise majority vote on remaining labels; (2) S-Ours-MV: Predict missing labels using annotator-specific models trained on the full training set, then aggregate all (remaining + padded) labels; Similarly, S-PADL-MV represents the best-performing baseline PADL (Liao et al., 2024) results. The results in Table 6 show that our method consistently outperforms baselines, validating that longitudinal annotator-wise modeling enhances aggregation reliability.
This section provides empirical evidence for our paradigm shift, further clarifying the motivation of this paper while depicting its practical utility value for the real world.
@@ 7.3. Video-specific Architecture
Different from the image input pipeline in the main paper, Figure 6 presents the video input pipeline specifically designed for the AMER, etc. video datasets. Specifically, given a video input $V\in\mathbb{R}^{T\times H\times W\times 3}$, a frozen pre-trained image encoder (Sun et al., 2023) first extracts frame features, which are then further compressed by the image Q-Former (Li et al., 2023) using a specific number of compression queries, typically 32, to alleviate subsequent computational costs (Li et al., 2023). The frame position embedding is then added and input to the video Q-Former. Subsequently, we assign a learnable query token in a Q-Former for modeling each annotator $A_{k}$ ($k=1,\dots,n$), then all annotator-specific queries in cross-attention simultaneously interact with input features to capture their diverse tendencies and obtain the annotator-specific feature. Finally, the specific annotator features are mapped through a fully connected layer to an appropriate feature dimension and then connected to each annotator’s corresponding classifier to output classification results.
For qualitative understanding, the focus regions of the video input are based on all frames. Different annotators’ varying levels of focus on different frames indicate their behavior pattern differences: annotator 1 shows more focus on the final frames, while annotator n focuses more on the middle frames. This dataset is for video emotion classification. Analyzing the annotation and visualization results, we see that annotators labeled emotions “anger” and “worried”. The differences in the frame segments they focused on might partially contribute to the variations in their judgments.